OpenStudy (anonymous):

Solve for x,y,z by elimination: 1)x-y+z=-1 x+y+3z=-3 2x-y+2z=0 2)x-y-2z=4 -x+2y+z=1 -x+y-3z=11

OpenStudy (anonymous):

1. x=4, y=2, z=-3 2. x=0, y=2, z=-3

OpenStudy (anonymous):

How did you get there? Can you show me the process please?

OpenStudy (anonymous):

x-y+z=-1 x+y+3z=-3 ---------------add 2x+4z=-4 x+y+3z=-3 2x-y+2z=0 --------------add 3x+5z=-3 (2x+4z=-4)*3 (3x+5z=-3)*2 6x+13z=-12 6x+10z=-6 -------------subtract 3z=-6 z=-3 3x+5z=-3 3x+5(-3)=-3 3x=-3+15 x= 12/3=4 x-y+z=-1 4-y-3=-1 y=4-3+1=2 you do the 2nd

OpenStudy (anonymous):

So how do you know when to add or subtract or is it in that precise order (add, add, subtract)?